OpenStudy (anonymous):

how to factor this binomial? 25x^2+1

OpenStudy (anonymous):

(-i+5 x) (i+5 x)

OpenStudy (asnaseer):

\[a^2+b^2=(a+ib)(a-ib)\]

OpenStudy (mertsj):

That is the sum of two squares and it is prime within the set of real numbers.

OpenStudy (mertsj):

What class are you taking?

OpenStudy (anonymous):

algebra...

OpenStudy (mertsj):

So have you had imaginary numbers yet?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

Then put down that this binomial cannot be factored and is prime.

OpenStudy (mertsj):

If it were the difference of two squares, it would be factorable but since it is the sum of two squares, it is prime.

OpenStudy (anonymous):

so what's the difference of sum/diff of squares?

OpenStudy (mertsj):

x^2 is a square 16 is a square x^+16 is the sum of two squares and cannot be factored x^2-16 is the difference of two squares and can be factored.

OpenStudy (mertsj):

Sum means add. Difference means subtract.

OpenStudy (anonymous):

so whenever there is a - sign it can be factored and when + is prime?

OpenStudy (mertsj):

Yes. If those signs join two squares.

OpenStudy (asnaseer):

Mertsj: why do say this must be prime? (e.g. when x=1, we get \(25x^2+1=26\) which is not a prime number.

OpenStudy (mertsj):

Prime means has no factors except itself and 1

OpenStudy (asnaseer):

so mean the "expression" is prime rather than the number that it produces?

OpenStudy (mertsj):

Yes

OpenStudy (asnaseer):

ok - thx for clarifying

OpenStudy (mertsj):

5 is prime

OpenStudy (mertsj):

x^2+4 is prime

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

got you!!!

OpenStudy (mertsj):

Good

Directrix (directrix):

x^2 + 4 is a prime expression over the set of Reals but is not prime over the set of Complex numbers.